OpenStudy (anonymous):

(7,2) (8_6) find the slope

OpenStudy (anonymous):

to find slope m = (y2-y1)/(x2-x1) m =slope suppose (7,2) is (x1,y1) and (8,6) is (x2,y2) then plug those into the formula m=(6-2)/(8-7) = 4

OpenStudy (anonymous):

answer correct